# Multicriteria Optimization and Decision Making

### v1

@@ 1.5 Formal Definition of Pareto Dominance
The theory of ordered sets is an essential analytical tool in multi-objective optimization and decision analysis. Different types of orders can be defined by means of axioms on binary relations, and, if we restrict ourselves to vector spaces, also geometrically.
Next we will first show how different types of orders are defined as binary relations that satisfy a certain axioms [^1] Using here the term ’axiom’ to refer to an elementary statement that is used to define a class of objects (as promoted by, for instance, Rudolf Carnap[16]) rather than viewing them as self-evident laws that do not require proof (Euclid’s classical view).. Moreover, we will highlight the essential differences between common families of ordered sets, such as preorders, partial orders, linear orders, and cone orders.
The structure of this chapter is as follows: After reviewing the basic concept of binary relations, we define some axiomatic properties of pre-ordered sets, a very general type of ordered sets. Then we define partial orders and linear orders as special type of pre-orders. The difference between linear orders and partial orders sheds a new light on the concept of incomparability and the difference between multicriteria and single criterion optimization. Later, we discuss techniques how to visualize finite ordered sets in a compact way, by so called Hasse diagrams. The remainder of this chapter deals with an alternative way of defining orders on vector spaces: Here we define orders by means of cones. This definition leads also to an intuitive way of how to visualize orders based on the concept of Pareto domination.
@@ 4.4 Inequality Constraints
Assume the objective and all constraint functions are convex in some $\epsilon$-neighborhood of ${\bf x}^{*}$, If ${\bf x}^{*}$ is a local minimum, then there exiss $\lambda_{1},\dots,\lambda_{m+1}$ such that KKT conditions are fulfilled.
@@
Note that if ${\bf x}^{*}$ is in the interior of the feasible region (a Slater point), all $g_{i}({\bf x})<0$ and thus $\lambda_{1}>0$.
@@ 4.5 Multiple Objectives
Karush Kuhn Tucker sufficient conditions for a solution to be Pareto optimal:
Let ${\bf x}^{*}$ be a feasible point. Assume that all objective functions are locally convex and all constraint functions are locally concave, and the Fritz John conditions hold in ${\bf x}^{*}$, then ${\bf x}^{*}$ is a local efficient point.
@@
In the unconstrained case we get the simple condition:
@@
In 2-dimensional spaces this criterion reduces to the observation, that either one of the objectives has a zero gradient (neccesary condition for ideal points) or the gradients are collinear as depicted in Fig. 4.4. A detailed description of the conditions for multiobjective optimization is given in [65].
@@ 5.3 Multi-Attribute Utility Theory
(Error converting to EPS. Consult Lieuwe)
Figure 5.6: In the left figure we see and examples for one-sided desirability function with parameters $y^{min}=-1,y^{max}=1,l\in\{0.5,1,1.5\}$. The left side displays a plot of two sided desirability functions for parameters $y^{min}=-1.y^{max}=1,$ $l\in\{0.5,1.0,1.5\}$, and $r$ being set to the same value than $l$.

### v2

@@ 1.5 Formal Definition of Pareto Dominance
The theory of ordered sets is an essential analytical tool in multi-objective optimization and decision analysis. Different types of order relations can be defined by means of axioms on binary relations and, if we restrict ourselves to vector spaces, also geometrically. Next, we will first show how different types of orders are defined as binary relations that satisfy a certain axioms [^1] Using here the term ’axiom’ to refer to an elementary statement that is used to define a class of objects (as promoted by, for instance, Rudolf Carnap[16]) rather than viewing them as self-evident laws that do not require proof (Euclid’s classical view).. Moreover, we will highlight the essential differences between common families of ordered sets, such as preorders, partial orders, linear orders, and cone orders.
The structure of this chapter is as follows. After reviewing the basic concept of binary relations, we define some axiomatic properties of pre-ordered sets, a very general type of ordered sets. Then we define partial orders and linear orders as a special type of pre-orders. The difference between linear orders and partial orders sheds new light on the concept of incomparability and the difference between multi-criteria and single-criterion optimization. Later, we discuss techniques how to visualize finite ordered sets in a compact way, by so-called Hasse diagrams. The remainder of this chapter deals with an alternative way of defining orders on vector spaces: Here we define orders by means of cones. This definition leads also to an intuitive way of how to visualize orders based on the concept of Pareto domination.
@@ 4.4 Inequality Constraints
Assume the objective and all constraint functions are convex in some $\epsilon$-neighborhood of ${\bf x}^{*}$, If ${\bf x}^{*}$ is a local minimum, then there exist $\lambda_{1},\dots,\lambda_{m+1}$ such that KKT conditions are fulfilled.
@@
Note that if ${\bf x}^{*}$ is in the interior of the feasible region (a Slater point), all $g_{i}({\bf x})<0$ and thus $\lambda_{1}>0$. This follows directly from the so-called non-slackness conditions in equation 4.14. Furthermore, let us observe that this equation makes sure that only active constraints in ${\bf x}^{*}$ enter the equation 4.13. Equation 4.13 essentially states that the negative gradient of the objective function must be contained in the polyhedral cone generated by the gradients of the active constraint’s gradients at ${\bf x}^{*}$. See, Section 1.5 for an introduction of the concept of a polyhedral cone.
@@ 4.5 Multiple Objectives
Karush Kuhn Tucker sufficient conditions for a solution to be Pareto optimal: Let ${\bf x}^{*}$ be a feasible point. Assume that all objective functions are locally convex and all constraint functions are locally concave, and the Fritz John conditions hold in ${\bf x}^{*}$, then ${\bf x}^{*}$ is a local efficient point.
@@
A simple way to understand these condition is to view the multipliers $\lambda_{i}$ as weights of a weighted sum scalarization with non-negative weights for the objective functions $\sum_{i=1}^{n}\lambda_{i}f_{i}(\mathbf{x})$, which we aim to minimize. In the unconstrained case we get the simple condition:
@@
In 2-dimensional spaces this criterion reduces to the observation, that either one of the objectives has a zero gradient (neccesary condition for ideal points) or the gradients are collinear as depicted in Fig. 4.4. We may in this special case use the angle between the gradients of the objective function as an indicator of the closeness to a locally efficient point, as it is for instance done in more recently proposed visualization methods for the landscape of multimodal multi-objective optimization problems (so-called gradient heat maps) [56].
This brief overview of the KKT condition for multiobjective optimization lacks rigorous proofs and instead relies on plausibility arguments. A more detailed mathematical treatment would necessitate introducing concepts that extend beyond the scope of this document. For a comprehensive derivation and detailed explanation of the conditions for multiobjective optimization, we suggest referring to [67].
@@ 5.3 Multi-Attribute Utility Theory
Figure 5.6: In the top figure we see and examples for one-sided desirability function with parameters $y^{min}=-1,y^{max}=1,l\in\{0.5,1,1.5\}$. The bottom figure displays a plot of two sided desirability functions for parameters $y^{min}=-1.y^{max}=1,$ $l\in\{0.5,1.0,1.5\}$, and $r$ being set to the same value than $l$.